OpenStudy (anonymous):

What is the exact area of a circle with a diameter of 18 feet? A. 36π square feet B. 81π square feet C. 324π square feet D. 1296π square feet

OpenStudy (acxbox22):

A=pi x r^2 d=2r 18=2r radius=9 

OpenStudy (anonymous):

the are = pi * r^2

OpenStudy (anonymous):

*area

OpenStudy (anonymous):

I got...36 is that right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Diameter is 18 so the radius =D/2=18/2=9 the area = pi * 9^2=81 pi

OpenStudy (anonymous):

I know its 254 but thats not in my answer choices~

OpenStudy (anonymous):

I get it now! than you!!!

OpenStudy (anonymous):

*thank

OpenStudy (anonymous):

YW